Please subscribe to the ACO-status mailing list for future ACO operational advisories and alerts.

The BSP2 dislodged from the bottom on 4/11/18, 00:53 UTC, moved to 25-30 mab and is floating with the current. CTD3, BPR pressure and fluorometer data may be affected.

New ACO instruments were installed during the ACO-3 cruise on 2014/11/01. Click here for ACO-3 videos.

Latest Seismogram Plots

Full Catalog of Seismogram Data Files

(Click on Thumbnails to Enlarge)


0.01-0.09 Hz Signal Envelope

4.5-19 Hz Signal Envelope

Latest seismogram time-series obtained with an OAS E-2PD hydrophone located on the JBOX frame of the observatory (Hydrophone Experiment Module). The sensor is nominally 1 m above the bottom. These graphs show low-frequency acoustic signals received by sources such as earthquakes, ships, and baleen whales.


0.01-0.09 Hz Signal Envelope

4.5-19 Hz Signal Envelope

0.01-0.09 Hz Signal Envelope

4.5-19 Hz Signal Envelope

0.01-0.09 Hz Signal Envelope

4.5-19 Hz Signal Envelope

0.01-0.09 Hz Signal Envelope

4.5-19 Hz Signal Envelope

0.01-0.09 Hz Signal Envelope

4.5-19 Hz Signal Envelope